# Retinex-RAWMamba: Bridging Demosaicing and Denoising for Low-Light RAW Image Enhancement

### v4

@@ Abstract
Low-light image enhancement, particularly in cross-domain tasks such as mapping from the raw domain to the sRGB domain, remains a significant challenge. Many deep learning-based methods have been developed to address this issue and have shown promising results in recent years. However, single-stage methods, which attempt to unify the complex mapping across both domains, leading to limited denoising performance. In contrast, two-stage approaches typically decompose a raw image with color filter arrays (CFA) into a four-channel RGGB format before feeding it into a neural network. However, this strategy overlooks the critical role of demosaicing within the Image Signal Processing (ISP) pipeline, leading to color distortions under varying lighting conditions, especially in low-light scenarios. To address these issues, we design a novel Mamba scanning mechanism, called RAWMamba, to effectively handle raw images with different CFAs. Furthermore, we present a Retinex Decomposition Module (RDM) grounded in Retinex prior, which decouples illumination from reflectance to facilitate more effective denoising and automatic non-linear exposure correction. By bridging demosaicing and denoising, better raw image enhancement is achieved. Experimental evaluations conducted on public datasets SID and MCR demonstrate that our proposed RAWMamba achieves state-of-the-art performance on cross-domain mapping. The code is available at https://github.com/Cynicarlos/RetinexRawMamba.
1 Introduction
@@
Demosaicing algorithms play a crucial role in converting RAW image to sRGB, with most traditional methods relying on proximity interpolation. Although some researchers have explored CNN-based approaches [42, 7] to map noisy RAW images to clean sRGB outputs, the limited receptive field inherent in convolutional networks often hampers their effectiveness in demosaicing tasks. To address this, Vision Transformers (ViTs) have been employed to expand the receptive field, but the attention mechanisms in ViTs are computationally intensive. The introduction of Mamba provides a more efficient balance between these trade-offs. However, existing Mamba scanning mechanisms do not adequately address the diverse characteristics of RAW images with different Color Filter Arrays (CFAs), highlighting the need for Mamba scanning methods specifically tailored to various CFAs.
@@
Additionally, previous methods [5, 17] for processing short-exposure RAW images often rely on a simple linear multiplication of a prior for exposure correction. Specifically, short-exposure RAW images, which contain significant noise, are multiplied by the exposure time ratio of the corresponding long-exposure image. This approach assumes uniform exposure across the image, which is often unrealistic and can result in sub-optimal denoising and inaccurate brightness. By leveraging the success of the Retinex theory in low-light enhancement tasks for RGB images [21, 25, 33], we introduce a Retinex-based dual-domain auxiliary exposure correction method, namely Retinex Decomposition Module (RDM), which decouples illumination and reflection and realize automatic nonlinear exposure correction to achieve more efficient denoising effect and more accurate brightness correction. Furthermore, given the significant differences in noise distribution between different RAW domain and sRGB domain, we build upon the idea of decoupling the task into two sub-tasks: denoising on the raw domain [43, 18, 35] and cross-domain mapping [30, 46, 34, 16].
In general, we propose a Retinex-based decoupling network (Retinex-RAWMamba) for RAW domain denoising and low-light enhancement shown in Fig. 2. Our method decouples the tasks of denoising and demosaicing into two distinct sub-tasks, effectively mapping noisy RAW images to clean sRGB images. Specifically, for demosaicing sub-task, we introduce RAWMamba to fully consider all pixels in the immediate neighborhood of a certain pixel by utilizing eight direction mechanism. For the denoising sub-task, we propose the Retinex Decomposition Module, which enhances both denoising performance and brightness correction. Additionally, we introduce a dual-domain encoding stage enhance branch designed to leverage the meticulously preserved detail features from the raw domain, thereby compensating for the information loss that occurs during the denoising phase.
@@
2 Related Work
2.1 Low Light Enhancement on Raw Domain
@@
In the Raw domain low-light enhancement task, researchers have proposed some innovative approaches. Since the task can be split into two sub-tasks, RAW domain denoising and color correction, some of the work only focuses on one of the sub-tasks. For example, on the raw domain denoising task, there are noise modeling with deep learning methods [2, 4, 8, 9, 39, 51], which ultimately compute evaluation metrics on the raw domain. After the release of the SID public dataset by Chen at al.[5] 2018, researchers have proposed many works that address both tasks simultaneously. These works can be further categorized into single-stage approaches and multi-stage approaches. Single-stage methods [5, 45] aims to map noisy raw to clean sRGB by training a single model. For instance, SID [5] only used a simple UNet to accomplish this task. DID [26] proposed a deep neural network based on residual learning for end-to-end extreme low-light image denoising. SGN [12] introduced a self-guided network, which adopted a top-down self-guidance architecture to better exploit image multi-scale information. Since the ISP undergoes many nonlinear transformations, it is still difficult to learn for a single neural network, and it can only be realized by piling up a large number of parameters, which leads to inefficiencies, and thus multi-stage methods came into being. Multi-stage methods [15, 42, 50] achieve better results by decoupling the tasks, this idea effectively reduces the ambiguity between different domains. For instance, Huang et al. [15] proposed intermediate supervision on the raw domain, while Dong et al. [7] did that on the monochrome domain. DNF [17] introduced a decoupled two-stage net with weight weight-shared encoder to reduce the number of parameters while achieving good results.
2.2 Mamba in Vision Task
@@
State Space Models (SSM) are recently introduced to deep learning since they can effectively model long range dependencies. For instance, [11] proposes a Structured State-Space Sequence (S4) model and recently, [10] proposes Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In addition to Mamba’s great work on NLP tasks, researchers have also made many attempts and achieved good results on visual tasks, such as classification [6, 44], segmentation [24, 37, 41, 29, 22], generation [31, 14], and image restoration [13, 32, 48, 49, 27, 1]. EfficientVMamba [27] presents the Efficient 2D Scanning (ES2D) method, utilizing atrous sampling of patches on the feature map to speed up training. VMamba [23] incorporates a Cross-Scan Module (CSM), which converts the input image into sequences of patches along the horizontal and vertical axes, and it enables the scanning of sequences in four distinct directions. That is, each pixel integrates information from the four surrounding pixels. FreqMamba [48] introduces complementary triple interaction structures including spatial Mamba, frequency band Mamba, and Fourier global modeling, which utilizes the complementarity between Mamba and frequency analysis for image deraining.
3 Method
3.1 Preliminaries
3.1.1 State Space Model (SSM)
@@
Finally, the models compute output through a global convolution as following:
@@
3.2 Overall Pipeline
The overall pipeline is shown in Fig. 2. First, we preprocess the low-exposure noisy single-channel raw image by multiplying it with the exposure time ratio of the long-exposure ground truth (GT). Then, based on the Color Filter Array (CFA) pattern, we pack it into a multi-channels input. Specifically, for Bayer format, we pack the input $\textbf{X}\in\mathbb{R}^{H\times W\times 1}$ into four channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$; for XTrans format, we pack the input into nine channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{3}\times\frac{W}{3}\times 9}$. Both stages of Retinex-RAWMamba are built upon the UNet-based [28] encoder-decoder architecture. The first stage of the overall framework is dedicated to raw domain denoising. Initially, The Retinex Decomposition Module (RDM) processes the input to generate two feature maps L and R, L will be multiplied by the original input to obtain $\textbf{X}_{in}$ and R will be used later. $\textbf{X}_{in}$ will pass the denoising stage and generate the first output $\textbf{O}_{1}\in\mathbb{R}^{H\times W\times C_{in}}$. Then $\textbf{X}_{in}$ will pass the demosaicing stage to generate the second output $\textbf{O}_{2}\in\mathbb{R}^{H\times W\times 3}$. The overall loss function is then calculated against both ground truth RAW and ground truth RGB images, providing the supervision signal for both domains and guiding the optimization of whole model.
@@ 3.3 RAWMamba
The detail of the our proposed scan mechanism is shown in Fig. 3 (c). Specifically, for a feature map $\textbf{{F}}\in{\mathbb{R}}^{C\times H\times W}$, we first flip its even rows (fer) and columns (fec) to get $\textbf{{F}}_{fer}\in{\mathbb{R}}^{C\times H\times W}$ and $\textbf{{F}}_{fec}\in{\mathbb{R}}^{C\times H\times W}$, respectively. Then we flatten $\textbf{{F}}_{fer}$ and $\textbf{{F}}_{fec}$ to get the first two directions’ scanning $\textbf{{F}}_{1}\in{\mathbb{R}}^{C\times HW}$ and $\textbf{{F}}_{2}\in{\mathbb{R}}^{C\times HW}$. Then we can get the feature of the oblique scan $\textbf{{F}}_{3}$ as following:
$(r,c)=(ceil(\frac{H\times W}{W+1}),(W+1))$ (5)
$\textbf{{F}}_{pad}=pad(flatten(\textbf{{F}}),(C,r\times c))$
$\textbf{{F}}_{reshaped}=reshape(\textbf{{F}}_{pad},(C,r,c))$
$\textbf{{F}}_{select}=ms(trans(\textbf{{F}}_{reshaped}),trans(\textbf{{mask}}))$
$\textbf{{F}}_{3}=reshape(\textbf{{F}}_{select},(C,H\times W))$
where $pad(\textbf{{X}},shape)$ is a function to pad X to new shape, mask and $\textbf{{F}}_{reshaped}$ have the same shape, and the first $H\times W$ elements of mask are true and the rest are false, trans(X) is a function to transpose $\textbf{{X}}\in{\mathbb{R}}^{C\times H\times W}$ to $\textbf{{X}}_{trans}\in{\mathbb{R}}^{C\times W\times H}$, $ms(\textbf{{X}},\textbf{{mask}})$ is a function to select elements from X based on the position where mask is true. The other oblique scan feature $\textbf{{F}}_{4}$ can be obtained similarly, and then we can invert these four features to get another four directions features, eight in total, which is $\{\textbf{{F}}_{i}\in{\mathbb{R}}^{C\times HW},i=1,2,...8\}$. At this point, the scanning of the eight directions is completed. And after the SSM, we get $\{\overline{\textbf{{F}}}_{i}\in{\mathbb{R}}^{C\times HW},i=1,2,...8\}$, we then merge them by adding them up and reshape these eight features to the original shape to get a single feature, that is,
@@
3.4 Retinex Decomposition
Low-light enhancement methods based on retinex theory have been successful in RGB domain [38, 40, 3], so we propose dual-domain Retinex Decomposition Module. RDM can decompose image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ into the reflection component $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$ and the illumination component $\textbf{{L}}\in{\mathbb{R}}^{H\times W\times C_{in}}$. The details of RDM are shown in Fig. 2 (a). The module first takes the average value of the input image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ in the channel dimension to obtain $\textbf{{M}}\in{\mathbb{R}}^{H\times W\times 1}$, concatenates it in the channel dimension, and then passes several convolutions and a GELU activate function to obtain the first output $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$, and then passes a $1\times 1$ convolution to obtain the light map L , which will be multiplied by the original input to pre-adjust the light. Specifically,

### v5

@@ Abstract
Low-light image enhancement, particularly in cross-domain tasks such as mapping from the raw domain to the sRGB domain, remains a significant challenge. Many deep learning-based methods have been developed to address this issue and have shown promising results in recent years. However, single-stage methods, which attempt to unify the complex mapping across both domains, leading to limited denoising performance. In contrast, existing two-stage approaches typically overlook the characteristic of demosaicing within the Image Signal Processing (ISP) pipeline, leading to color distortions under varying lighting conditions, especially in low-light scenarios. To address these issues, we propose a novel Mamba-based method customized for low light RAW images, called RAWMamba, to effectively handle raw images with different CFAs. Furthermore, we introduce a Retinex Decomposition Module (RDM) grounded in Retinex prior, which decouples illumination from reflectance to facilitate more effective denoising and automatic non-linear exposure correction, reducing the effect of manual linear illumination enhancement. By bridging demosaicing and denoising, better enhancement for low light RAW images is achieved. Experimental evaluations conducted on public datasets SID and MCR demonstrate that our proposed RAWMamba achieves state-of-the-art performance on cross-domain mapping. The code is available at https://github.com/Cynicarlos/RetinexRawMamba.
Index Terms:
RAW Image, Low Light, Mamba, ISP
publicationid: pubid: Copyright ©20xx IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
I Introduction
@@
Demosaicing algorithms play a crucial role in converting RAW image to sRGB, with most traditional methods relying on proximity interpolation. Although some researchers have explored CNN-based approaches [4, 5] to map noisy RAW images to clean sRGB outputs, the limited receptive field inherent in convolutional networks often hampers their effectiveness in demosaicing tasks. To address this, Vision Transformers (ViTs) have been employed to expand the receptive field, but the attention mechanisms in ViTs are computationally intensive. The introduction of Mamba provides a more efficient balance between these trade-offs [6, 7, 8]. However, existing Mamba scanning mechanisms do not adequately address the diverse characteristics of RAW images with different Color Filter Arrays (CFAs), highlighting the need for Mamba scanning methods specifically tailored to various CFAs.
@@
Additionally, previous methods [9, 10] for processing short-exposure RAW images often rely on a simple linear multiplication of a prior for exposure correction. Specifically, short-exposure RAW images, which contain significant noise, are multiplied by the exposure time ratio of the corresponding long-exposure image. This approach assumes uniform exposure across the image, which is often unrealistic and can result in sub-optimal denoising and inaccurate brightness. By leveraging the success of the Retinex theory in low-light enhancement tasks for RGB images [11, 12, 13], we introduce a Retinex-based dual-domain auxiliary exposure correction method, namely Retinex Decomposition Module (RDM), which decouples illumination and reflection and realize automatic nonlinear exposure correction. At the same time, we efficiently fuse the generated priors based on multi-scale fusion strategy[14, 15, 16, 17, 18, 19] to achieve more efficient denoising effect and more accurate brightness correction. Furthermore, given the significant differences in noise distribution between different RAW domain and sRGB domain, we build upon the idea of decoupling the task into two sub-tasks: denoising on the RAW domain and cross-domain mapping.
In general, we propose a Retinex-based decoupling network (Retinex-RAWMamba) for RAW domain denoising and low-light enhancement shown in Fig. 2. Our method decouples the tasks of denoising and demosaicing into two distinct sub-tasks, effectively mapping noisy RAW images to clean sRGB images. Specifically, for demosaicing sub-task, we introduce RAWMamba to fully consider all pixels in the immediate neighborhood of a certain pixel by utilizing eight direction mechanism. For the denoising sub-task, we propose the Retinex Decomposition Module, which enhances both denoising performance and brightness correction. Additionally, we introduce a dual-domain encoding stage enhance branch designed to leverage the meticulously preserved detail features from the raw domain, thereby compensating for the information loss that occurs during the denoising phase. Finally, We observe that the Gated Fusion Module (GFM) used in prior works led to unstable training within our framework. To address this, Domain Adaptive Fuse (DAF)is proposed to perform adaptive feature fusion with better stability and efficiency.
@@
II Related Work
II-A Low Light Enhancement on Raw Domain
@@
Raw image consists much more details than its corresponding RGB image, and it would be better to enhance the low light image on Raw domain than RGB domain, like some homomorphic filtering methods [20, 21, 22], which are a great inspiration for us. For the Raw domain low-light enhancement task, researchers have proposed some innovative approaches. Since the task can be split into two sub-tasks, RAW denoising and cross domain mapping. For example, on the raw domain denoising task, there are noise modeling with deep learning methods [23, 24, 25, 26, 27, 28], which ultimately compute evaluation metrics on the raw domain. After the release of the SID public dataset by Chen at al.[9] 2018, researchers have proposed many works that address both tasks simultaneously. These works can be further categorized into single-stage approaches and multi-stage approaches. Single-stage methods [9, 29] aims to map noisy raw to clean sRGB by training a single model. For instance, SID [9] only used a simple UNet to accomplish this task. DID [1] proposed a deep neural network based on residual learning for end-to-end extreme low-light image denoising. SGN [2] introduced a self-guided network, which adopted a top-down self-guidance architecture to better exploit image multi-scale information.
Since the ISP undergoes many nonlinear transformations, it is still difficult to learn for a single neural network, and it can only be realized by piling up a large number of parameters, which leads to inefficiencies, and thus multi-stage methods came into being. Multi-stage methods [30, 4, 31] achieve better results by decoupling the tasks, this idea effectively reduces the ambiguity between different domains. For instance, Huang et al. [30] proposed intermediate supervision on the raw domain, while Dong et al. [5] did that on the monochrome domain. DNF [10] introduced a decoupled two-stage net with weight-shared encoder to reduce the number of parameters while achieving good results. However, the weight-sharing module used across both domains may introduce cross-domain ambiguity, resulting in suboptimal performance.
II-B Deep Learning for ISP
The motivation for replacing hardware-based ISP systems with deep learning solutions stems from their superior capability in reconstructing lost image information while mitigating cumulative processing errors inherent in traditional multi-stage ISP pipelines [32]. Recent advancements in this field have demonstrated significant progress through diverse methodological innovations [33, 34]. CycleISP framework proposed by Zamir et al. [35] features a bidirectional architecture containing complementary RGB2RAW and RAW2RGB conversion branches, enhanced by an adaptive color correction module to simulate camera imaging pipelines and generate paired data for dual-domain denoising. For demosaicing optimization, Xu et al. [36] developed a hierarchical processing architecture called DemosaicFormer, implementing coarse reconstruction followed by pixel-level refinement. Addressing cross-device adaptability, Perevozchikov et al. [37] pioneered an unpaired learning paradigm for RAW-to-RAW translation across heterogeneous camera sensors, enabling flexible deployment of neural ISPs on unseen devices. MetaISP framework by Souza et al. [38] introduced metadata-aware domain adaptation, leveraging EXIF parameters and illuminant estimation to achieve cross-device characteristic translation. In computational efficiency optimization, Guan et al. [39] innovated a grouped deformable convolution mechanism for joint denoising and demosaicing, strategically allocating independent offset parameters across kernel groups to balance accuracy and latency. Collectively, these advancements validate the viability of deep learning-based ISP solutions through comprehensive technical explorations spanning data synthesis, architectural design, and deployment optimization.
II-C Mamba in Vision Task
@@
State Space Models (SSM) are recently introduced to deep learning since they can effectively model long range dependencies. For instance, [40] proposes a Structured State-Space Sequence (S4) model and recently, [41] proposes Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In addition to Mamba’s great work on NLP tasks, researchers have also made many attempts and achieved good results on visual tasks[42, 43], such as classification [44, 45], segmentation [46, 47, 48, 49, 50, 51], anomaly detection[52], point cloud learning[53], generation [54, 55], and image restoration [8, 7, 56, 57, 58].
EfficientVMamba [57] presents the Efficient 2D Scanning (ES2D) method, utilizing atrous sampling of patches on the feature map to speed up training. VMamba [6] incorporates a Cross-Scan Module (CSM), which converts the input image into sequences of patches along the horizontal and vertical axes, and it enables the scanning of sequences in four distinct directions. That is, each pixel integrates information from the four surrounding pixels. VMambaIR [7] proposes omni selective scan mechanism to overcome the unidirectional modeling limitation of SSMs by efficiently modeling image information flows in all six directions for RGB image images, which added two more directions along the channel dimension. In contrast, we combine the characteristics of AI ISP task, proposing the eight direction scanning mechanism and Retinex decomposition module to overcome the uneven lighting of the low light RAW images. FreqMamba [59] introduces complementary triple interaction structures including spatial Mamba, frequency band Mamba, and Fourier global modeling, which utilizes the complementarity between Mamba and frequency analysis for image deraining. Similarly, WalMaFa[60] proposes a novel Wavelet-based Mamba with Fourier adjustment model. RetinexMamba[58] directly integrates Mamba into RetinexFormer[61] for low light RGB image enhancement without any changes about the Mamba itself. Li et al.[62] combines contrastive learning and Mamba to achieve semi-supervised learning. However, most of the existing scanning mechanism in the mentioned Mamba have limitations. One direction scanning in the original Mamba [41] that is used for sequence prediction task does usually not do well in vision tasks, since the image has two dimensions and each pixel are usually related to its surrounding pixels instead of only its next one as the sequence. Therefore, most vision Mamba adopt the four directions scanning mechanism. Li et al. [63] utilizes four directions scanning strategy for underwater image enhancement. Nevertheless, it just consider the up, down, left and right pixels and ignore the spatial continuity, which is not the best approach in our task. Therefore, RetinexRawMamba is proposed to address this issue by considering eight scanning directions. And to address the uneven light of the low light RAW image, we introduced the retinex decomposition module to estimate the light components to adaptively correct the light and color by multi-scale fusion, which can also complement mamba’s ability to capture local information.
III Method
III-A Preliminaries
III-A1 State Space Model (SSM)
@@
Finally, the models compute output through a global convolution as follows:
@@
III-B Overall Pipeline
The overall pipeline is shown in Fig. 2. First, we preprocess the low-exposure noisy single-channel raw image by multiplying it with the exposure time ratio of the long-exposure ground truth (GT). Then, based on the Color Filter Array (CFA) pattern, we pack it into a multi-channels input. Specifically, for Bayer format, we pack the input $\textbf{X}\in\mathbb{R}^{H\times W\times 1}$ into four channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$; for XTrans format, we pack the input into nine channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{3}\times\frac{W}{3}\times 9}$. Both stages of Retinex-RAWMamba are built upon the UNet-based [64] encoder-decoder architecture. While the U-Net architecture remains a parsimonious choice for dense prediction tasks, our research introduces a domain-specific architectural innovation that transcends conventional designs: the first-ever integration of Mamba with Retinex-guided multi-scale prior fusion for low-light RAW image enhancement. This approach addresses a critical research gap—existing Mamba applications in 2024 have overlooked RAW data’s unique challenges, as validated by our comprehensive literature survey (Section II.C). Furthermore, The experimental results (Section IV.B) also prove that this architecture design has a significant performance improvement over the baseline architecture.
The first stage of the overall framework is dedicated to raw domain denoising. Initially, The Retinex Decomposition Module (RDM) processes the input to generate two feature maps L and R, L will be multiplied by the original input to obtain $\textbf{X}_{in}$ and R will be used later. $\textbf{X}_{in}$ will pass the denoising stage and generate the first output $\textbf{O}_{1}\in\mathbb{R}^{H\times W\times C_{in}}$. Then $\textbf{X}_{in}$ will pass the demosaicing stage to generate the second output $\textbf{O}_{2}\in\mathbb{R}^{H\times W\times 3}$. The overall loss function is then calculated against both ground truth RAW and ground truth RGB images, providing the supervision signal for both domains and guiding the optimization of whole model.
@@ III-C RAWMamba
The detail of the our proposed scan mechanism is shown in Fig. 3 (c). We will firstly obtain eight directions scanning features, which is $\{\textbf{{F}}_{i}\in{\mathbb{R}}^{C\times HW},i=1,2,...8\}$. At this point, the scanning of the eight directions is completed. And after the SSM, we get $\{\overline{\textbf{{F}}}_{i}\in{\mathbb{R}}^{C\times HW},i=1,2,...8\}$, we then merge them by adding them up and reshape these eight features to the original shape to get a single feature, that is,
@@
III-D Retinex Decomposition and Dual-domain Encoding Stage Enhance Branch
Low-light enhancement methods based on retinex theory have been successful in RGB domain [65, 66, 61], so we propose dual-domain Retinex Decomposition Module. Our RDM is inspired by RetinexFormer[61], which just use several convolutions to estimate the illumination map and the reflect map. The previous retinex-based methods are one stage network and the prior features are used in both encoding and decoding stage, but we remove two decoding stages prior fusion in our two stages network to reduce the computational complexity. RDM can decompose image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ into the reflection component $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$ and the illumination component $\textbf{{L}}\in{\mathbb{R}}^{H\times W\times C_{in}}$. The details of RDM are shown in Fig. 2 (a). The module first takes the average value of the input image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ in the channel dimension to obtain $\textbf{{M}}\in{\mathbb{R}}^{H\times W\times 1}$, concatenates it in the channel dimension, and then passes several convolutions and a GELU activate function to obtain the first output $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$, and then passes a $1\times 1$ convolution to obtain the light map L , which will be multiplied by the original input to pre-adjust the light. Specifically,